OpenStudy (anonymous):

how do u write 2x+3y=0 in slope intercept form?

OpenStudy (anonymous):

So in y=mx+b No, b is here. 3y=-2x y= -(2/3)x

OpenStudy (anonymous):

how do u write x=y+5 and 2x-2y=12 slope intercept form??

OpenStudy (anonymous):

First one is: -y=-x+5 ---> y = x-5 Second one is: 2x-2y=12 x-y=6 -y=-x+6 y=x-6

OpenStudy (anonymous):

solve by addition tell whether system is independent dependent or inconsistent, -3x+y=3 2x-3y=5

OpenStudy (anonymous):

Try it yourself first please.